# Rethinking Prompt Optimizers: From Prompt Merits to Optimization

### v2

@@ D.3 BBH Experimental Results
date_understanding 57.77 36.4 35.6 49.2 52.4 34 31.2 50 64.8
disambiguation_qa 66.4 38.4 37.2 48.8 54.8 44.8 40.8 68.8 71.6
hyperbaton 81.6 54.4 50.4 50.4 54.4 56.8 48.4 62 90.8
logical_deduction_five_objects 58.8 27.2 24.4 47.2 42.4 30 23.6 52 66.8
logical_deduction_seven_objects 49.2 29.6 20.4 39.6 36.8 21.2 18.4 51.6 63.2
logical_deduction_three_objects 82 36.8 51.2 42.8 39.2 40.4 34.8 65.6 80.8
movie_recommendation 63.2 38.4 26.8 38.4 29.2 32 27.2 34 77.6
penguins_in_a_table 56.85 28.77 25.34 38.36 54.79 31.51 25.34 51.37 71.23
reasoning_about_colored_objects 56.8 35.6 38 23.6 41.6 26 18.4 58.8 63.2
ruin_names 38 27.6 41.6 42 28.8 45.6 28.4 73.6 59.2
salient_translation_error_detection 43.2 36.8 38.8 28.8 31.2 39.2 23.6 54.8 72
snarks 67.42 53.37 51.12 49.44 66.85 60.11 46.07 83.15 76.97
word_sorting 95.2 92.4 90 92 86.4 84 3.2 100 100
temporal_sequences 38 37.6 39.2 33.2 42 29.2 28.8 66.8 84.4
tracking_shuffled_objects_five_objects 31.2 39.6 33.2 35.2 32.8 20.4 23.2 21.2 28
tracking_shuffled_objects_seven_objects 15.2 30 26.4 16.4 28.4 18.8 15.2 18 30.4
tracking_shuffled_objects_three_objects 33.6 32.4 33.2 33.2 64.8 31.6 31.6 29.6 31.6
causal_judgement 49.2 56.68 51.87 51.79 56.15 62.57 52.41 59.36 80.21
formal_fallacies 51.2 57.2 54.8 53.2 52.8 53.2 53.2 59.2 83.2
navigate 61.6 58.8 44.4 44.8 48.4 47.2 42 60 42.4
web_of_lies 51.2 55.2 51.6 51.6 59.2 51.2 49.6 49.6 71.6
sports_understanding 77.6 67.2 54.8 77.2 61.2 58 46 49.2 75.6
boolean_expressions 48.4 61.2 54 54.4 54 52.8 55.6 36.4 66.4
multistep_arithmetic_two 45.6 14.4 14.4 41.2 16 57.6 18.4 18 84.4
object_counting 39.6 35.2 60.4 55.2 34 60.8 49.6 53.6 86.8

### v3

@@ D.1 BBH Experimental Results
Gemma2
-9b
Qwen3
-8b
DeepSeek
-Qwen3
DeepSeek
-LLaMA3
date_understanding 44.8 30.8 28.8 31.6 42 25.6 27.2 62 60.8 51.6 58 32.8 60.4
disambiguation_qa 63.2 32.8 31.2 32.4 30 36.4 33.6 68.8 65.6 67.6 67.6 56 56
hyperbaton 75.6 48.4 48.4 48.4 48.4 51.6 48.4 78 87.6 85.2 90.8 86.4 58
logical_deduction_five_objects 49.2 22.4 18 28 33.2 24.4 19.6 50 58.8 62.8 62.8 35.2 40
logical_deduction_seven_objects 42.4 17.6 14.8 30 27.2 30.4 15.2 42.8 54.4 65 57.2 31.2 31.2
logical_deduction_three_objects 66.4 32.8 31.6 33.6 36.4 35.2 31.6 71.2 75.6 51.6 86 64.8 67.6
movie_recommendation 48 26.8 22.4 33.2 23.6 23.2 22.4 52.8 76.4 66.4 63.6 61.6 45.6
penguins_in_a_table 54.11 21.23 22.6 31.51 46.58 26.71 23.97 50.68 67.12 57.53 54.79 31.51 57.53
reasoning_about_colored_objects 47.6 32.4 32 16.4 29.2 30 16.4 46 51.6 30 47.6 37.2 46.4
ruin_names 37.2 24 30.8 35.2 27.2 41.6 28.4 66.8 51.2 45.6 74 53.6 39.6
salient_translation_error_detection 32 26.8 23.2 21.6 28 34 16.8 52 71.6 66.4 66.4 55.6 27.6
snarks 60.11 44.94 46.07 46.07 50.56 57.3 46.07 81.46 74.16 67.56 68.54 58.43 47.75
word_sorting 88 82.4 80 84 80 52 1.2 100 100 90 100 100 100
temporal_sequences 61.2 19.2 28.4 28.4 33.2 28 28.4 62 83.6 52.8 77.2 55.2 40
tracking_shuffled_objects_five_objects 28.4 19.2 20 20.8 30.8 20 20 17.2 40.4 30 36.8 15.2 46
tracking_shuffled_objects_seven_objects 13.2 24 14.4 13.6 22 11.2 12.8 13.6 18.4 28.8 30.4 14 24
tracking_shuffled_objects_three_objects 27.2 29.2 24.4 26.8 49.2 30 30.8 28 31.2 45.6 54.4 30.4 30
causal_judgement 62.57 49.2 51.87 51.34 49.2 47.59 51.87 53.48 79.68 87.17 67.91 58.82 49.2
formal_fallacies 54 52.8 53.2 53.6 52.4 51.6 53.2 52 77.2 54.4 63.6 49.6 55.2
navigate 57.6 53.6 42 41.6 41.6 41.2 42 52.8 57.2 81.2 81.2 42 42
web_of_lies 46.4 48.8 48.8 49.2 54 44.4 46.4 46.8 68.4 68.4 50 45.6 47.2
sports_understanding 70 50 46.4 70.4 54 53.6 43.2 45.6 66.8 68.4 69.6 80.8 46
boolean_expressions 34.8 56.8 52.4 54.8 51.2 48 54 30.4 62.8 64.8 100 24.8 54
multistep_arithmetic_two 33.6 12.4 9.6 36 14 46 16 16.8 85.6 75.2 96.4 58.8 86
object_counting 36 33.2 32.8 50.4 33.2 48.8 46.8 49.6 90 71.2 76.8 58.8 66.4
- Qwen2-7b Tulu2-7b-dpo LLaMA2-7b-chat-hf LLaMA3-8b-instruct Gemma2-9b Qwen3-8b
date_understanding 55.6 36 34 48 60 58
disambiguation_qa 66 37.2 35.2 35.2 69.2 70.4
hyperbaton 78.4 50 48.4 49.6 86.8 93.6
logical_deduction_five_objects 55.6 25.2 19.2 34.8 73.8 73.8
logical_deduction_seven_objects 44 28.4 21.6 38.8 67.2 52.4
logical_deduction_three_objects 75.2 34.4 42.4 40.4 58 76.8
movie_recommendation 53.2 31.2 24.4 37.6 75.6 76.4
penguins_in_a_table 56.16 27.4 23.29 36.99 62.53 53.42
reasoning_about_colored_objects 50.8 34 35.2 20.8 30.8 58.8
ruin_names 38 28.4 40 39.6 46 73.2
salient_translation_error_detection 38.4 34 37.2 18.8 73.34 73.34
snarks 62.36 50 50.56 48.88 72.81 71.35
word_sorting 94.4 87.2 87.6 90 100 100
temporal_sequences 64.4 36.8 36.8 31.2 58.4 78.8
tracking_shuffled_objects_five_objects 29.6 38.4 34.8 33.2 32.8 40.4
tracking_shuffled_objects_seven_objects 14.4 24.8 23.2 14.4 25.2 30.4
tracking_shuffled_objects_three_objects 30 31.6 31.6 31.2 56.4 56.4
causal_judgement 55.08 52.41 51.87 52.41 92.57 75.82
formal_fallacies 54 54 54 52 65.2 70.2
navigate 57.2 58.8 43.2 42.4 88.8 88.4
web_of_lies 52.4 49.6 49.2 51.6 57.2 52.8
sports_understanding 76 64.8 49.2 79.6 78.8 84
boolean_expressions 41.6 60.4 53.6 50 72 100
multistep_arithmetic_two 41.2 12.6 12 41.6 84.4 93.2
object_counting 39.2 36.4 58.4 50.8 68 68.8
Qwen2
-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
Gemma2
-9b
Qwen3
-8b
DeepSeek
-Qwen3
DeepSeek
-LLaMA3
date_understanding 56.8 33.2 35.2 48.8 29.6 29.2 57.6 57.6 37.2 61.2
disambiguation_qa 63.2 34 34.4 40.4 44.4 34.4 64.8 64.8 66.4 58
hyperbaton 80 53.2 50 49.6 51.6 48.8 85.6 90.8 84.4 61.6
logical_deduction_five_objects 52.8 25.2 24 31.2 29.2 20.4 76.8 57.2 40.8 37.6
logical_deduction_seven_objects 46.4 28.8 19.2 35.6 24.8 15.2 70 56 37.2 30.8
logical_deduction_three_objects 78.4 32.8 45.6 39.2 34.8 32 54.8 83.2 72.4 80.4
movie_recommendation 61.2 22.8 24.4 37.6 23.2 26.8 75.6 72.8 62.4 54.4
penguins_in_a_table 54.11 23.97 23.29 36.3 29.45 23.97 60.85 56.16 26.71 73.29
reasoning_about_colored_objects 52.4 34 35.6 20.8 28.4 18 30.4 51.6 24.4 40
ruin_names 38.4 28 36.8 35.6 38 28.4 46 72 48.4 35.2
salient_translation_error_detection 38.4 27.6 35.2 21.2 33.6 18.8 71.2 72.4 42.8 17.6
snarks 66.85 53.37 53.37 48.88 53.93 46.07 70 71.91 57.87 43.82
word_sorting 92.4 90 87.6 90 72 3.2 84 100 100 100
temporal_sequences 66.4 26.4 38.4 30.8 35.6 28.4 58 79.2 60.4 30.8
tracking_shuffled_objects_five_objects 29.6 29.2 31.2 34 20 19.6 31.2 40.4 15.6 17.2
tracking_shuffled_objects_seven_objects 14.4 28 23.2 15.2 14 13.2 28.8 30.8 14 15.6
tracking_shuffled_objects_three_objects 27.6 31.6 31.6 31.6 31.6 32 56.8 56.8 33.6 32.8
causal_judgement 53.48 50.8 52.04 52.94 52.94 51.87 79.14 74.33 48.13 56.68
formal_fallacies 52 53.2 54.14 54 52.4 53.2 62.8 69.2 53.2 53.2
navigate 59.6 58.4 44.8 48.4 44.8 42 85.6 85.6 43.2 42
web_of_lies 49.2 51.6 48.8 49.6 49.2 48.8 68.4 54 48.4 49.2
sports_understanding 76.4 61.2 53.2 74.8 55.2 45.6 75.6 80.8 84.4 46
boolean_expressions 45.6 63.6 55.6 53.2 50.4 54 64 100 37.6 53.6
multistep_arithmetic_two 36.8 10 17.6 37.2 46.8 16.8 80.4 96.4 63.6 89.6
object_counting 34.4 32.4 34.8 50 54.4 43.6 69.2 76.8 74 70.8
Qwen2-
7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
Gemma2
-9b
Qwen3
-8b
DeepSeek
-Qwen3
DeepSeek
-LLaMA3
date_understanding 55.2 38.4 29.2 36 30 27.2 58.4 57.6 56.4 58.8
disambiguation_qa 61.6 34.8 31.2 40.4 32 33.2 63.6 63.6 53.6 56.4
hyperbaton 78.8 33.2 48.4 60 52 48.4 94 90.8 74.8 63.2
logical_deduction_five_objects 50 21.6 22.8 34.4 23.6 20.4 73.6 57.2 43.2 50.8
logical_deduction_seven_objects 46.4 25.2 20.4 39.2 23.2 15.6 72.8 56 39.6 36.4
logical_deduction_three_objects 70.8 30.4 42 39.2 38.8 32 55.2 83.2 80.8 78
movie_recommendation 62 36.8 23.97 31.6 34.4 24.4 72.8 72.8 54.4 42.8
penguins_in_a_table 54.11 36.3 23.29 30.82 52.05 23.29 71.92 56.16 48.63 73.97
reasoning_about_colored_objects 50.8 31.6 33.2 24.8 25.6 16.4 40 51.6 31.6 32
ruin_names 46 34.4 36 38.8 40 29.2 46 72 49.2 40
salient_translation_error_detection 37.6 32.4 33.6 24.4 37.2 18.8 72.4 72.4 42.4 34.8
snarks 62.92 46.07 50.56 46.07 55.62 46.07 72.81 71.91 55.06 45.51
word_sorting 92.4 92.4 86.4 92 70 4.8 86.4 100 100 100
temporal_sequences 63.2 27.6 28.4 28.8 40 28.8 57.2 79.2 56.8 46.4
tracking_shuffled_objects_five_objects 31.2 24.8 32.4 36.8 20.8 20.8 27.6 40.4 19.2 16.8
tracking_shuffled_objects_seven_objects 16.8 25.6 24.8 14.4 14.4 16.4 24 30.8 13.2 12.8
tracking_shuffled_objects_three_objects 31.2 30.4 31.6 30.4 31.6 30.8 42 54.4 32.4 31.2
causal_judgement 62.57 53.48 51.87 51.34 55.08 52.41 83.42 74.33 61.5 57.22
formal_fallacies 53.2 54.4 54 52.4 53.2 53.2 65.2 69.2 56.4 53.2
navigate 59.2 60.8 42 46.4 53.2 42 74 85.6 43.6 42
web_of_lies 58.8 48 53.2 45.2 49.2 48.8 92 54 41.2 42
sports_understanding 81.2 66.4 47.26 71.6 55.6 46 80.8 96.4 71.6 57.6
boolean_expressions 39.6 58.8 53.6 61.2 39.6 54.4 74 100 38.8 52.4
multistep_arithmetic_two 16.4 11.6 17.2 39.6 24.4 16.4 83.2 96.4 25.6 31.6
object_counting 36.4 33.6 61.2 50.8 55.2 43.2 69.2 69.2 49.6 63.6
Qwen2
-7b
Tulu2-7b
-dpo
LLaMA2-7b
-chat-hf
LLaMA3-8b
-instruct
LLaMA2-13b
-chat-hf
LLaMA3.2-3b
-instruct
LLaMA3.2-1b
-instruct
Tulu2-70b
-dpo
LLaMA3.3-70b
-instruct
Gemma2
-9b
Qwen3
-8b
DeepSeek
-Qwen3
DeepSeek
-LLaMA3
date_understanding 57.77 36.4 35.6 49.2 52.4 34 31.2 50 64.8 62.8 56.8 49.2 63.6
disambiguation_qa 66.4 38.4 37.2 48.8 54.8 44.8 40.8 68.8 71.6 70.4 69.2 65.6 60.8
hyperbaton 81.6 54.4 50.4 50.4 54.4 56.8 48.4 62 90.8 87.6 91.2 87.2 62.8
logical_deduction_five_objects 58.8 27.2 24.4 47.2 42.4 30 23.6 52 66.8 77.8 77.8 43.2 45.6
logical_deduction_seven_objects 49.2 29.6 20.4 39.6 36.8 21.2 18.4 51.6 63.2 70.8 57.2 40 33.2
logical_deduction_three_objects 82 36.8 51.2 42.8 39.2 40.4 34.8 65.6 80.8 65.2 74.4 75.6 87.6
movie_recommendation 63.2 38.4 26.8 38.4 29.2 32 27.2 34 77.6 77.6 75.6 66 54.4
penguins_in_a_table 56.85 28.77 25.34 38.36 54.79 31.51 25.34 51.37 71.23 62.27 57.53 52.05 76.03
reasoning_about_colored_objects 56.8 35.6 38 23.6 41.6 26 18.4 58.8 63.2 40 45.2 18 54.8
ruin_names 38 27.6 41.6 42 28.8 45.6 28.4 73.6 59.2 48 70.8 68.4 40
salient_translation_error_detection 43.2 36.8 38.8 28.8 31.2 39.2 23.6 54.8 72 74.4 74.4 49.6 30.4
snarks 67.42 53.37 51.12 49.44 66.85 60.11 46.07 83.15 76.97 73.56 74.16 74.72 56.18
word_sorting 95.2 92.4 90 92 86.4 84 3.2 100 100 100 100 100 100
temporal_sequences 38 37.6 39.2 33.2 42 29.2 28.8 66.8 84.4 58.8 79.2 61.6 29.6
tracking_shuffled_objects_five_objects 31.2 39.6 33.2 35.2 32.8 20.4 23.2 21.2 28 33.2 40.4 24.8 15.2
tracking_shuffled_objects_seven_objects 15.2 30 26.4 16.4 28.4 18.8 15.2 18 30.4 29.2 30.4 15.6 13.6
tracking_shuffled_objects_three_objects 33.6 32.4 33.2 33.2 64.8 31.6 31.6 29.6 31.6 56.8 56.8 34 31.6
causal_judgement 49.2 56.68 51.87 51.79 56.15 62.57 52.41 59.36 80.21 95.24 78.07 64.17 54.55
formal_fallacies 51.2 57.2 54.8 53.2 52.8 53.2 53.2 59.2 83.2 66.4 76 56 53.2
navigate 61.6 58.8 44.4 44.8 48.4 47.2 42 60 42.4 88.4 88.8 42.4 42
web_of_lies 51.2 55.2 51.6 51.6 59.2 51.2 49.6 49.6 71.6 80 50.4 55.6 52.8
sports_understanding 77.6 67.2 54.8 77.2 61.2 58 46 49.2 75.6 86 84.4 84.4 45.6
boolean_expressions 48.4 61.2 54 54.4 54 52.8 55.6 36.4 66.4 77.6 100 42.8 60.4
multistep_arithmetic_two 45.6 14.4 14.4 41.2 16 57.6 18.4 18 84.4 85.6 96.4 78 90
object_counting 39.6 35.2 60.4 55.2 34 60.8 49.6 53.6 86.8 69.2 72.8 75.6 74.4